OpenStudy (anonymous):

Simplify. square root of 3 times square root of 21

OpenStudy (bibby):

\(\sqrt{a} \sqrt{b}=\sqrt{a*b}\)

OpenStudy (anonymous):

\[ sqrt{3} * sqrt{21} = sqrt{63}=sqrt{9*7} = 3sqrt{7} \]